# Sparse identification of quasipotentials via a combined data-driven method

### v1

@@ 2.3 Sparse symbolic regression of the quasipotential function
Based on the neural network approximations for the potential $V$ and circulatory vector $\mathbf{g}$, we further identify a sparse symbolic expression for the orthogonal decomposition of $\mathbf{f}$ that facilitates physical interpretation of the system. This is achieved by projecting the dynamics over a minimal set of functions from a candidate library $\bm{\Theta}(\mathbf{X})=[\theta_{1}(\mathbf{X}),\dots,\theta_{q}(\mathbf{X})]$, which is appropriately selected standing for possible functions. For $V$ and $\mathbf{g}$, the symbolic representation reads
@@
where $\bm{\Xi}_{v}$ and $\bm{\Xi}_{g}$ are the vectors containing the unknown coefficients to be determined. In addition, $\mathbf{X}$ represents a $NM$-by-$d$ matrix containing the data points $\{X_{i}^{j,0}\}$ from the sampled trajectories with multiple initial conditions. The domain of the symbolic regression does not necessarily coincides with that of the neural networks. Indeed, it is convenient not to overextend the identification far from the metastable states of the system. Outer boundaries of $V$ are inherently unreliable due to scarcity of data. We limit the selection over $\mathbf{X}$ by defining a subset of $\{X_{i}^{j,0}\}$ below a potential threshold value $\tau$, i.e. $V_{\theta}(\mathbf{x})<\tau$.
@@
Now, existing sparse regression approaches aim to directly match observable data or byproducts such as their derivatives [29]. Unfortunately, directly inferring a symbolic regression of the quasipotential from data is not feasible. The essential information about the energy landscape remains concealed unless we possess prior knowledge on how to divide the dynamics into downhill and circulatory components. Therefore, we structure the symbolic identification to extract the quasipotential function from the decomposition provided by the neural networks. Specifically, a target matrix for the identification is set as
@@
where $R(\cdot)$ is a $l_{0}$-regularization term that promotes sparsity and minimizes over-fitting and W is an auxiliary variable which is introduced here to enable relaxation and partial minimization in order to improve the conditioning of the problem and tackle the non-convexity of the optimization [30]. In addition, $\lambda$ and $\nu$ are hyper-parameters that control the strength of regularization and relaxation, respectively. As indicated from Eq. (12),
@@
We solve the constrained optimization problem (14), (16) by performing an initialization step for the coefficient matrix $\bm{\Xi}$ (see A), obtaining $V^{\text{Symb}}$. The symbolic expression for the Freidlin-Wentzell quasipotential is its scalar multiple, $U^{\text{Symb}}(\mathbf{x})=2\cdot V^{\text{Symb}}(\mathbf{x})$.
@@ 3 Results
To provide a context for our study and to validate the efficacy of our machine learning approach, we investigate a system with a precisely known quasipotential (Sec. 3.1). Subsequently, we employ the combined data-driven technique to analyze the dynamics of a nanomechanical graphene resonator (Sec. 3.2).
In both the applications, we create a synthetic set of data points by sampling trajectories of the system and taking snapshots of the trajectories. In the first part of the method (Sec. 2.2), we train neural networks to reconstruct an orthogonal decomposition of the vector field. Details for sampling the data and training the neural networks are provided in B. In the second part of the method (Sec. 2.3), we identify sparse symbolic expressions for the downhill and circulatory components that define the dynamics of the system.
@@ 3.1 Example 1: An archetypal model with exact quasipotential
$\left\{\begin{array}[]{l}\displaystyle\dot{x}=-2\left(x^{3}-x\right)-y-z\vspace{1mm}\\ \displaystyle\dot{y}=-y+2\left(x^{3}-x\right)\vspace{1mm}\\ \displaystyle\dot{z}=-z+2\left(x^{3}-x\right)\end{array}\right.$ (17)
The system has two stable equilibrium points at $A=(-1,0,0)$ and $B=(1,0,0)$ and the quasipotential with respect to the two states is given by $U(\mathbf{x})=x^{4}-2x^{2}+y^{2}+z^{2}+1$.
We generate $5000$ trajectories of the system by numerically simulating the dynamics (17) with initial states sampled from the computational domain $\Omega=[-2,2]\times[-1.5,1.5]\times[-1.5,1.5]$. Fig. 1 (a) shows a plot of five generated trajectories. A dataset $\{X_{i}^{j,0},X_{i}^{j,h}\}$ is then constructed by taking snapshots from the generated trajectories. From the data, we first reconstruct an orthogonal decomposition of the vector field by training the neural networks $V_{\theta}(\mathbf{x})$ and $\mathbf{g}_{\theta}(\mathbf{x})$ with the loss function (7). Subsequently, we perform the sparse symbolic regression for the downhill and circulatory components in the decomposition. The identified components are showcased in Fig. 1 (b) and (c). Specifically, we take the basis functions for the regression as polynomials in $(x,y,z)$ up to fifth order for both $V^{\text{Symb}}$, $\mathbf{g}^{\text{Symb}}$. The target matrix is constructed with the neural network solutions over a subset of $\{X_{i}^{j,0}\}$ built with a potential threshold $\tau=\min_{\Omega}V_{\theta}(\mathbf{x})+2$. In the learning problem (14), we take the parameters $\lambda=0.1$ and $\nu=10^{-5}$.
@@
Figure 1: Combined data-driven method applied to the system (17) with $5000$ sampled trajectories. (a) Five generated trajectories. (b) The identified downhill component $-\nabla V^{\text{Symb}}(x,y,z)$. (c) The identified circulatory component $\mathbf{g}^{\text{Symb}}(x,y,z)$. (d) The exact quasipotential $U(x,y,z)$. (e) The identified quasipotential $U^{\text{Symb}}(x,y,z)$. All plots are projected on the $(x,y)$-plane. In panel (b)-(c), the line thickness shows the flow velocity.
@@
$d\mathbf{x}_{t}=\mathbf{f}(\mathbf{x}_{t})dt+\sqrt{\epsilon}\ dW_{t},\quad t>0$ (19)
where $\epsilon$ is a parameter controlling the strength of the noise and $W_{t}$ is a Wiener process. When $\epsilon$ is small, the invariant distribution of the system (19) can be approximated via the quasipotential $U(\mathbf{x})$,
@@
where $Z$ is a normalization constant [12].
For the system (19), computing the invariant distribution $p_{\epsilon}(\mathbf{x})$ with the identified quasipotential $U^{\text{Symb}}(\mathbf{x})$ allows us to accurately estimate the normalization constant $Z$. This calculation is detailed in C. The estimated constants, by accounting for various values of $\epsilon$ ($0.05$, $0.1$ and $0.2$), are
@@ 3.2 Example 2: Nanomechanical graphene resonator
$\left\{\begin{array}[]{l}\displaystyle\dot{P}=\dfrac{{\omega_{0}}^{2}-{\omega_{F}}^{2}}{2\omega_{F}}Q-\zeta P+\dfrac{3}{8}\dfrac{\alpha}{\omega_{F}}Q\left(P^{2}+Q^{2}\right)\vspace{2mm}\\ \displaystyle\dot{Q}=-\dfrac{{\omega_{0}}^{2}-{\omega_{F}}^{2}}{2\omega_{F}}P-\zeta Q-\dfrac{3}{8}\dfrac{\alpha}{\omega_{F}}P\left(P^{2}+Q^{2}\right)-\dfrac{\beta}{2\omega_{F}}\text{.}\end{array}\right.$ (23)
@@
describe the slow dynamics of a graphene membrane in terms of the $P$ and $Q$ variables that are the slowly varying in-phase and out-of-phase components of the motion, and obtained by applying a rotating wave approximation to the actual fast dynamics [6]. We take the following parameters in the dynamical equation: $\omega_{0}\!=\!1$, $\omega_{F}\!=\!1.0018$, $\zeta\!=\!0.00045$ and $\alpha\!=\!33$, and $\beta\!=\!1.4\!\times\!10^{-5}$. These describe a graphene drum in the classical duffing bistable region where both high and low amplitude oscillations occurs. The dynamics has two stable fix points at $A\approx(-0.007,-0.011)$ and $B\approx(0.004,-0.001)$, as shown in Fig. 3 (a).
@@
Figure 3: (a) Five generated trajectories. (b) The observed dataset $\{X_{i}^{j,0}\}$, where the solid closed curve indicates the boundary of the region $\{\mathbf{x}\in\Omega:V_{\theta}(\mathbf{x})<\tau\}$ with a threshold value $\tau=\min_{\Omega}V_{\theta}(\mathbf{x})+6\times 10^{-9}$. (c) The dataset for the data matrix $\mathbf{X}$ sampled using the confined region (Right).
@@
We create a synthetic dataset $\{X_{i}^{j,0},X_{i}^{j,h}\}$ by generating $2000$ trajectories of the system (23) with initial conditions in a domain $\Omega=[-0.02,0.02]\times[-0.02,0.02]$. Five generated trajectories are shown in Fig. 3 (a). From the data, we learn an orthogonal decomposition of the vector field by training the neural networks $V_{\theta}(P,Q)$ and $\mathbf{g}_{\theta}(P,Q)$ with the loss function (7). Subsequently, we identify symbolic expressions for the components in the decomposition. A target matrix $\mathbf{G}(\mathbf{X})$ is constructed with the neural network solutions over a subset of $\{X_{i}^{j,0}\}$ built with a potential threshold $\tau=\min_{\Omega}V_{\theta}(\mathbf{x})+6\times 10^{-9}$. The full dataset used for training the neural networks, along with the reduced dataset utilized during the regression phase, are displayed in Fig. 3 (b) and (c). The basis functions in the identification are taken as polynomials in $(P,Q)$ up to fourth and third orders for $V^{\text{Symb}}$ and $\mathbf{g}^{\text{Symb}}$, respectively. We set the parameters as $\lambda=10^{-9}$ and $\nu=10^{-2}$ in the problem (14). An initial guess for the coefficient matrix $\bm{\Xi}_{v}$ and $\bm{\Xi}_{g}$ is obtained by performing regression for $V^{\text{Symb}}(\mathbf{x})$ and $\mathbf{g}^{\text{Symb}}(\mathbf{x})$ separately (see A).
@@
Figure 4: Combined data-driven method applied to the system (23) with $2000$ sampled trajectories. (a) The identified downhill component $-\nabla V^{\text{Symb}}(P,Q)$. (b) The identified circulatory component $\mathbf{g}^{\text{Symb}}(P,Q)$. (c) The quasipotential $U(P,Q)$ estimated using a standard ordered upwind method [11]. (d) The potential $U_{\theta}=2V_{\theta}(P,Q)$ learned using the neural networks. (e) The potential $U^{\text{Symb}}(P,Q)$ identified using the sparse symbolic regression. In panel (a)-(b), the line thickness shows the flow velocity.
@@
$U^{\text{Symb}}(P,Q)$ $=(8.12\times 10^{-9})-(1.56\times 10^{-6})P+(3.58\times 10^{-7})Q$ (24)
$\quad+(9.65\times 10^{-6})P^{2}+(1.71\times 10^{-5})PQ+(5.83\times 10^{-4})Q^{2}$
$\quad+0.023P^{3}+0.039P^{2}Q+0.047PQ^{2}+0.098Q^{3}$
$\quad+1.05P^{4}+1.88P^{3}Q+3.13P^{2}Q^{2}+2.47PQ^{3}+4.08Q^{4}.$
@@
$g_{1}^{\text{symb}}(P,Q)$ $=-(5.15\times 10^{-7})-(4.58\times 10^{-4})P-(1.88\times 10^{-3})Q$ (25)
$\quad+0.027P^{2}+0.048PQ-0.005Q^{2}$
$\quad+2.01P^{3}+14.61P^{2}Q+4.17PQ^{2}+11.99Q^{3},$
$g_{2}^{\text{symb}}(P,Q)$ $=-(6.61\times 10^{-6})+(1.77\times 10^{-3})P-(3.47\times 10^{-5})Q$
$\quad+0.012P^{2}+0.034PQ+0.116Q^{2}$
$\quad-11.60P^{3}+2.69P^{2}Q-9.68PQ^{2}+6.92Q^{3}.$
In Fig. 4 (a)-(b), we show plots of the identified downhill and circulatory components $\nabla V^{\text{Symb}}(P,Q)$, $\mathbf{g}^{\text{Symb}}(P,Q)$ in the decomposition of the vector field. Unlike the example in Sec. 3.1, we do not have knowledge of the true quasipotential. It is important to emphasize that this is the most common situation in real-world applications, and it further underlines the need of a data-driven approach. Thus, we find an alternative way to compare our estimation. Fig. 4 (c) reports a reference solution for the quasipotential $U$ using a standard ordered upwind method with a mesh of $1000\times 1000$ grid points [2]. The mesh-based method is compared with the identification provided by the neural network (Fig. 4 (d)), and the sparse symbolic regression (Fig. 4 (e)). From the figure, one can observe that the neural network (panel (d)) and symbolic (panel (e)) solutions agree well with the reference solution (panel (f)), which indicate the capability of the proposed method to infer the quasipotential from the observed data.

### v2

@@ 2.3 Sparse symbolic regression of the quasipotential function
Based on the neural network approximations for the potential $V$ and circulatory vector $\mathbf{g}$, We determine a sparse symbolic expression for the orthogonal decomposition of $\mathbf{f}$, which facilitates its physical interpretation and enables analytical treatment. This is achieved by projecting the dynamics over a minimal set of functions from a candidate library $\bm{\Theta}(\mathbf{X})=[\theta_{1}(\mathbf{X}),\dots,\theta_{q}(\mathbf{X})]$, which is appropriately selected standing for possible functions. For $V$ and $\mathbf{g}$, the symbolic representation reads
@@
where $\bm{\Xi}_{v}$ and $\bm{\Xi}_{g}$ are the vectors of the unknown coefficients to be determined. In addition, $\mathbf{X}$ represents a NMNM-by-$d$ matrix containing the data points $\{X_{i}^{j,0}\}$ from the sampled trajectories with multiple initial conditions. The data domain of the symbolic regression does not necessarily coincides with that of the neural networks. Indeed, it is convenient not to overextend the identification far from the metastable states of the system. Outer boundaries of $V$ are inherently unreliable due to scarcity of data. We limit the selection over $\mathbf{X}$ by defining a subset of $\{X_{i}^{j,0}\}$ below a potential threshold value $\tau$, i.e. $V_{\theta}(\mathbf{x})<\tau$.
@@
Now, existing sparse regression approaches aim to directly match observable data or byproducts such as their derivatives [Quade2018]. Unfortunately, directly inferring a symbolic regression of the quasipotential from data is not feasible. The essential information about the energy landscape remains concealed unless we possess prior knowledge on how to divide the dynamics into downhill and circulatory components. Therefore, we structure the symbolic identification to extract the quasipotential function from the decomposition provided by the neural networks. Specifically, a target matrix for the identification is set as
@@
where $R(\cdot)$ is a $l_{0}$-regularization term that promotes sparsity and minimizes over-fitting and W is an auxiliary variable which is introduced here to enable relaxation and partial minimization in order to improve the conditioning of the problem and tackle the non-convexity of the optimization [Zheng2019]. In addition, $\lambda$ and $\nu$ are hyper-parameters that control the strength of regularization and relaxation, respectively. As indicated from Eq. (15),
@@
We note that, in addition to the constraint (19), it is possible to incorporate physical knowledge to guide the sparse regression toward a specific representation of the dynamics (see, e.g., [Lathourakis2024]). We solve the constrained optimization problem (17), (19) by performing an initialization step for the coefficient matrix $\bm{\Xi}$ (see Appendix A), obtaining $V^{\text{Symb}}$. The symbolic expression for the Freidlin-Wentzell quasipotential is its scalar multiple, $U^{\text{Symb}}(\mathbf{x})=2\cdot V^{\text{Symb}}(\mathbf{x})$. Finally, we note that projecting onto a set of continuous functions accurately handles only continuously differentiable energy surface topologies. However, the regression method can be extended to accommodate discontinuities, cf. [Mangan2019] and references therein.
@@ 3 Results
To provide a context for our study and test the efficacy of our machine learning approach, we investigate an archetypal system with known quasipotential (Sec. 3.1). This example is further examined by introducing noise and extended to four dimensions in Sec. 3.1.1 and Sec. 3.1.2, respectively. Subsequently, we employ the combined data-driven technique to analyze a system of engineering interest: the dynamics of a nanomechanical graphene resonator (Sec. 3.2).
In both the applications, we create a synthetic set of data points by sampling trajectories of the system and taking snapshots of the trajectories. In the first part of the method (Sec. 2.2), we train neural networks to reconstruct an orthogonal decomposition of the vector field. Details for sampling the data and training the neural networks are provided in Appendix B. In the second part of the method (Sec. 2.3), we identify sparse symbolic expressions for the downhill and circulatory components that define the dynamics of the system.
@@ 3.1 Example 1: An archetypal model with exact quasipotential
$\left\{\begin{array}[]{l}\displaystyle\dot{x}=-2\left(x^{3}-x\right)-y-z,\vskip 2.84526pt\\ \displaystyle\dot{y}=-y+2\left(x^{3}-x\right),\vskip 2.84526pt\\ \displaystyle\dot{z}=-z+2\left(x^{3}-x\right),\end{array}\right$. (20)
The system has two stable equilibrium points at A=(1,0,0)A=(-1,0,0) and $B=(1,0,0)$ and the quasipotential with respect to the two states is given by $U(\mathbf{x})=x^{4}-2x^{2}+y^{2}+z^{2}+1$.
We generate $5000$ trajectories of the system by numerically simulating the dynamics with initial states sampled from the computational domain $\Omega=[-2,2]\times[-1.5,1.5]\times[-1.5,1.5]$. Figure 1 (a) shows a plot of five generated trajectories. A dataset $\{X_{i}^{j,0},X_{i}^{j,h}\}$ is then constructed by taking snapshots from the generated trajectories. From the data, we first reconstruct an orthogonal decomposition of the vector field by training the neural networks $V_{\theta}(\mathbf{x})$ and $\mathbf{g}_{\theta}(\mathbf{x})$ with the loss function (9). Subsequently, we perform the sparse symbolic regression for the downhill and circulatory components in the decomposition. The identified components are showcased in Fig. 1 (b) and (c). Specifically, we take the basis functions for the regression as polynomials in $(x,y,z)$ up to fifth order for both $V^{\text{Symb}}$, $\mathbf{g}^{\text{Symb}}$. The target matrix is constructed with the neural network solutions over a subset of $\{X_{i}^{j,0}\}$ built with a potential threshold $\tau=\min_{\Omega}V_{\theta}(\mathbf{x})+2$. In the learning problem (17), we take the parameters $\lambda=0.1$ and $\nu=10^{-5}$.
@@
Figure 1: Combined data-driven method applied to the system (20) with $5000$ sampled trajectories. (a) Five generated trajectories. (b) The identified downhill component $\nabla V^{\text{Symb}}(x,y,z)$. (c) The identified circulatory component $\mathbf{g}^{\text{Symb}}(x,y,z)$. (d) The exact quasipotential $U(x,y,z)$. (e) The identified quasipotential $U^{\text{Symb}}(x,y,z)$. All plots are projected on the $(x,y)$-plane. In panels (b) and (c), the line thickness shows the flow velocity.
Figure 2: Distributions of the coefficients for each library term in the identified potential $U^{\text{Symb}}(\mathbf{x})$ in $50$ independent runs for the system in Eq. (20). Panel (a)-(e) indicate the coefficient distributions for the nontrivial terms ($1$, $x^{2}$, $y^{2}$, $z^{2}$ and $x^{4}$) as in $U(\mathbf{x})$, while panel (f) shows that learned coefficients are zero in all runs for any of the remaining library terms.
@@
We implement the proposed method in $50$ independent runs. On each run, we generate $5000$ trajectories from randomly sampled initial states, train the neural networks and perform the sparse symbolic regression, in the same way as described above. In Fig. 2, we show the distributions of the learned coefficients for each library term in $U^{\text{Symb}}(\mathbf{x})$. The relative root mean square error of the computed coefficients in $U^{\text{Symb}}(\mathbf{x})$ corresponding to the nontrivial terms ($1$, $x^{2}$, $y^{2}$, $z^{2}$ and $x^{4}$) is $8.87\times 10^{-4}\pm 3.48\times 10^{-4}$. The statistics of the results confirm that the proposed method is able to predict the coefficients of nontrivial terms with high accuracy and identify the trivial terms of the quasipotential (see Fig. 3).
Figure 3: Error of the symbolic model with respect the exact one for the 50 repetitions of the data-driven symbolic identification.
@@
$d\mathbf{x}_{t}=\mathbf{f}(\mathbf{x}_{t})dt+\sqrt{\epsilon}\ dW_{t},\quad t>0,$ (22)
where $\epsilon$ is a parameter controlling the strength of the noise and $W_{t}$ is a Wiener process. When $\epsilon$ is small, the invariant distribution of the system (22) can be approximated via the quasipotential $U(\mathbf{x})$ [Freidlin1998, lin2022computing, lin2023computing],
@@
where $Z$ is a normalization constant.
For the system (22), computing the invariant distribution $p_{\epsilon}(\mathbf{x})$ with the identified quasipotential $U^{\text{Symb}}(\mathbf{x})$ allows us to accurately estimate the normalization constant $Z$. This calculation is detailed in Appendix C. The estimated constants, by accounting for various values of $\epsilon$ ($0.05$, $0.1$ and $0.2$), are
@@
3.1.1 Application of the method to noisy data
A more practical situation is that the system is under the influence of random perturbations. To validate the ability of the method for dealing with such systems, we consider the system in Eq. (20) perturbed by small noise,
$\left\{\begin{array}[]{l}\displaystyle\dot{x}=-2\left(x^{3}-x\right)-y-z+\xi_{1}(t),\vskip 2.84526pt\\ \displaystyle\dot{y}=-y+2\left(x^{3}-x\right)+\xi_{2}(t),\vskip 2.84526pt\\ \displaystyle\dot{z}=-z+2\left(x^{3}-x\right)+\xi_{3}(t),\end{array}\right$. (26)
where $\xi(t)=(\xi_{1}(t),\xi_{2}(t),\xi_{3}(t))$ is a white noise with $\langle\xi_{j}(t)\xi_{k}(0)\rangle=2\epsilon\delta_{jk}\delta(t)$ in which $\epsilon$ is a parameter specifying the magnitude of the noise. In the example, we take the parameter $\epsilon=0.01$.
We generate a dataset of $5000$ noisy trajectories by numerically simulating the dynamics using the Euler-Maruyama scheme. A plot of the sampled trajectories, as compared against the ones sampled from the deterministic system (20) with the same initial conditions, is shown in Fig. 5 (a). Then, with the noisy trajectories, we implement the method for training the neural networks with the loss function (9) and performing the sparse symbolic regression by solving the problem (17), (19). All the parameters in the data generation, neural network training and symbolic regression are taken as the same ones in Sec. 3.1. The identified expressions of the quasipotential and circulatory component are
[…]
In Fig. 5 (b) and (c), we show plots of the predicted trajectories generated using the learned dynamics $\dot{\mathbf{x}}=\mathbf{f}^{\text{Symb}}(\mathbf{x})$ and the identified potential $U^{\text{Symb}}(\mathbf{x})$. The results demonstrate the capability of the proposed method for dealing with noisy trajectories.
Figure 5: Combined data-driven method applied to the system (26) with $5000$ sampled noisy trajectories. (a) Six generated trajectories and the ones sampled from the deterministic system (20) with the same initial conditions. (b) Six trajectories generated by simulating the learned dynamics $\dot{\mathbf{x}}=\mathbf{f}^{\text{Symb}}(\mathbf{x})$ and the ones sampled from the system (20) with the same initial states. (c) The identified quasipotential $U^{\text{Symb}}(x,y,z)$. All plots are projected on the $(x,y)$-plane.
3.1.2 Extension to a 4D system
We extend the three-dimensional non-gradient dynamical system of Eq. (20) to four dimensions by introducing an additional state variable $w$ and its corresponding equation $\dot{w}$, formulated in the same structural form as $\dot{y}$ and $\dot{z}$:
$\left\{\begin{array}[]{l}\displaystyle\dot{x}=-2\left(x^{3}-x\right)-y-z-w,\vskip 2.84526pt\\ \displaystyle\dot{y}=-y+2\left(x^{3}-x\right),\vskip 2.84526pt\\ \displaystyle\dot{z}=-z+2\left(x^{3}-x\right),\vskip 2.84526pt\\ \displaystyle\dot{w}=-w+2\left(x^{3}-x\right),\end{array}\right$. (27)
This example follows the same structure as the 3D case in Sec. 3.1, with two stable equilibrium points at A=(1,0,0,0)A=(-1,0,0,0) and $B=(1,0,0,0)$. The quasipotential consistent with the vector field is given by $U(\mathbf{x})=x^{4}-2x^{2}+y^{2}+z^{2}+w^{2}+1$. Using the same procedure as in the 3D case, we construct the decomposition into downhill and circulatory components and recover the symbolic form of the quasipotential. A comparison between the data-driven identification and the exact solution is reported in Table 2.
Exact expression     Identified expression
$U(\mathbf{x})$ $\quad x^{4}-2x^{2}+y^{2}+z^{2}+w^{2}+1$ $\hskip 14.22636pt0.976x^{4}-1.972x^{2}+0.951y^{2}$
$\hskip 0.0pt+1.007z^{2}+1.014w^{2}+0.994$
$\nabla V(\mathbf{x})$ $\begin{bmatrix}2x^{3}-2x\\ y\\ z\\ w\end{bmatrix}$ $\hskip 14.22636pt\begin{bmatrix}1.952x^{3}-1.972x\\ 0.951y\\ 1.007z\\ 1.014w\end{bmatrix}$
$\mathbf{g}(\mathbf{x})$ $\begin{bmatrix}-y-z-w\\ 2x^{3}-2x\\ 2x^{3}-2x\\ 2x^{3}-2x\end{bmatrix}$ $\hskip 14.22636pt\begin{bmatrix}-1.0y-1.014z-0.993w\\ 1.997x^{3}-1.990x\\ 2.002x^{3}-1.998x\\ 2.016x^{3}-2.015x\end{bmatrix}$
The comparison between the generated trajectories and the analytically reconstructed ones is shown in Fig. 6, illustrating the applicability of the method to quasipotential problems in dimensions higher than three, a topic that has not been thoroughly investigated in the literature.
Figure 6: Phase plots of six observed trajectories (Left) and the trajectories generated by simulating the dynamics $\dot{\mathbf{x}}=\mathbf{f}^{\text{Symb}}(\mathbf{x})$ with the same initial states (Right). The plots are projected on the $(x,y,z)$-space.
@@ 3.2 Example 2: Nanomechanical graphene resonator
$\left\{\begin{array}[]{l}\displaystyle\dot{P}=\dfrac{{\omega_{0}}^{2}-{\omega_{F}}^{2}}{2\omega_{F}}Q-\zeta P+\dfrac{3}{8}\dfrac{\alpha}{\omega_{F}}Q\left(P^{2}+Q^{2}\right)\vskip 5.69054pt\\ \displaystyle\dot{Q}=-\dfrac{{\omega_{0}}^{2}-{\omega_{F}}^{2}}{2\omega_{F}}P-\zeta Q-\dfrac{3}{8}\dfrac{\alpha}{\omega_{F}}P\left(P^{2}+Q^{2}\right)-\dfrac{\beta}{2\omega_{F}}\text{.}\end{array}\right$. (28)
@@
describe the slow dynamics of a graphene membrane in terms of the $P$ and $Q$ variables that are the slowly varying in-phase and out-of-phase components of the motion, and obtained by applying a rotating wave approximation to the actual fast dynamics [Dolleman2019]. We take the following parameters in the dynamical equation: resonant frequency $\omega_{0}\!=\!1$, frequency of the excitation $\omega_{F}\!=\!1.0018$, damping ratio $\zeta\!=\!0.00045$, cubic stiffness coefficient $\alpha\!=\!33$ and amplitude of the excitation $\beta\!=\!1.4\!\times\!10^{-5}$. These describe a graphene drum in the classical duffing bistable region where both high and low amplitude oscillations occur. The dynamics has two stable fix points at $A\approx(-0.007,-0.011)$ and $B\approx(0.004,-0.001)$, as shown in Fig. 7 (a).
@@
Figure 7: (a) Five generated trajectories of Eq. (28). (b) The observed dataset $\{X_{i}^{j,0}\}$, where the solid closed curve indicates the boundary of the region $\{\mathbf{x}\in\Omega:V_{\theta}(\mathbf{x})<\tau\}$ with a threshold value $\tau=\min_{\Omega}V_{\theta}(\mathbf{x})+6\times 10^{-9}$. (c) The dataset for the data matrix $\mathbf{X}$ sampled using the confined region.
@@
We create a synthetic dataset $\{X_{i}^{j,0},X_{i}^{j,h}\}$ by generating $2000$ trajectories of the system (28) with initial conditions in a domain $\Omega=[-0.02,0.02]\times[-0.02,0.02]$. From the data, we learn an orthogonal decomposition of the vector field by training the neural networks $V_{\theta}(P,Q)$ and $\mathbf{g}_{\theta}(P,Q)$. Subsequently, we identify symbolic expressions for the components in the decomposition. A target matrix $\mathbf{G}(\mathbf{X})$ is constructed with the neural network solutions over a subset of $\{X_{i}^{j,0}\}$ built with a potential threshold $\tau=\min_{\Omega}V_{\theta}(\mathbf{x})+6\times 10^{-9}$. The full dataset used for training the neural networks, along with the reduced dataset utilized during the regression phase, are displayed in Fig. 7 (b) and (c). The basis functions in the identification are taken as polynomials in $(P,Q)$ up to fourth and third orders for $V^{\text{Symb}}$ and $\mathbf{g}^{\text{Symb}}$, respectively. We set the parameters as $\lambda=10^{-9}$ and $\nu=10^{-2}$ in the problem (17). An initial guess for the coefficient matrix $\bm{\Xi}_{v}$ and $\bm{\Xi}_{g}$ is obtained by performing regression for $V^{\text{Symb}}(\mathbf{x})$ and $\mathbf{g}^{\text{Symb}}(\mathbf{x})$ separately, as we detail in Appendix A.
@@
Figure 8: Combined data-driven method applied to the system (28) with $2000$ sampled trajectories. (a) The identified downhill component $\nabla V^{\text{Symb}}(P,Q)$. (b) The identified circulatory component $\mathbf{g}^{\text{Symb}}(P,Q)$. (c) The quasipotential $U(P,Q)$ estimated using a standard ordered upwind method [Cameron2012]. (d) The potential $U_{\theta}=2V_{\theta}(P,Q)$ learned using the neural networks. (e) The potential $U^{\text{Symb}}(P,Q)$ identified using the sparse symbolic regression. In panels (a) and (b), the line thickness shows the flow velocity.
@@
[…]
@@
[…]
In Fig. 8 (a) and (b), we show plots of the identified downhill and circulatory components $\nabla V^{\text{Symb}}(P,Q)$, $\mathbf{g}^{\text{Symb}}(P,Q)$ in the decomposition of the vector field. Figure 8 (c) reports a reference solution for the quasipotential $U$ using a standard ordered upwind method with a mesh of $1000\times 1000$ grid points [Nolting2016]. The mesh-based method is compared with the identification provided by the neural network (Fig. 8 (d)), and the sparse symbolic regression (Fig. 8 (e)). From the comparison with the reference solution, one can observe that the neural network (panel (d)) and symbolic (panel (e)) capture the topology of the energy landscape well. However, we observe a misalignment between the outbound regions of $U^{\text{Symb}}$ (Fig. 8 (e)) and the reference solution. We attribute this imprecision to the low data density, as shown in Fig. 7 (b), and the high flow velocity, illustrated in Fig. 8 (a), both of which require special handling during the decomposition step. The error between the reference quasipotential estimated using the standard ordered upwind method (Fig. 5(c)) and the two data-driven steps of our method is presented in Appendix E. This highlights a limitation of the approach: it is fully data-driven and does not require prior knowledge of the system, but its performance depends heavily on the quantity and quality of the available data. We emphasize that the dataset can be improved through filtering, smoothing, and interpolation techniques to mitigate these issues. Additionally, the representation capacity of the polynomials in the candidate library for symbolic regression may be limited in capturing such variation. Despite this, the solution $U^{\text{Symb}}(P,Q)$ gives a physically interpretable form of the quasipotential. In this example, we also implement the data-driven method in $50$ independent runs and obtain similar potential $U^{\text{Symb}}$ as in Fig. 8 (e), where plots of the coefficient distribution for each library term in $U^{\text{Symb}}(P,Q)$ are shown in Fig. 10 of Appendix D.